# Entity-based Reinforcement Learning for Autonomous Cyber Defence

### v2

@@ 6.2 Generalisation to unseen topologies
The second set of experiments we conducted was aimed at evaluating how well entity-based agents generalise at test-time to network sizes not seen during training. As the baseline MLP policies do not support dynamically-sized inputs that are different to the ones used at train-time, they cannot be evaluated on varying network topologies at test-time.
To assess the generalisation ability of entity-based policies across different network sizes, we took three policies trained in the Random training regime, one for each network size of 10, 20 and 40. We then evaluated each of the policies on random networks for each of the network sizes. This means that we evaluated a policy trained on random 10-node networks in the same context of random 10-node networks but also on 20 and 40-node networks (‘out-of-training’ context). All the evaluations were carried out over $1000$ episodes, and so $1000$ different random networks.
Figure 3 displays the distribution of episodic rewards over 1,000 test episodes in the form of box plots, obtained from the evaluations of the three RogueNet agents trained on different network sizes. Subfigure 3(a) presents evaluations of agents trained on 10, 20 and 40-node networks respectively, and evaluated on random 10-node networks at test-time (eval_rand_k_on_10 with $\mathbf{k\in\{10,20,40\}}$). Subfigures 3(b) and 3(c) presents the corresponding evaluations on 20-node networks, 3(b), and 40-node networks, 3(c). Summary statistics of the same experiments are presented in Table 1.
Overall, we can observe a negligible difference in the performance of the three entity-based agents across the different network sizes when evaluated at test-time. For example, the agent trained on 40-node networks performs just as well when evaluated on 10-node networks as the agent that was trained natively on 10-node networks. In fact, it shows slightly lower variance in episodic reward than the agent trained on 10 nodes, with $0.87$ standard deviation against the $1.31$ standard deviation of the native 10-node agent. The only exception is the agent trained on 20-node networks and evaluated on 40-node networks, which shows slightly lower mean reward compared to the other two agents. As can be seen from the summary statistics results in Table 1, this case presents higher standard deviation in the rewards distribution (0.332) compared to the other two cases (0.187 and 0.161). These results suggest an ability for an entity-based agent to learn a general policy that performs consistently on any network size. The generality of these results would need to be verified in more complex environment settings or simulators, and different varieties of network topologies.
@@ 7 Discussion
In this paper, we have demonstrated the basic functionality and the strong generalisation capabilities of an entity-based RL method through the use of the Entity Gym framework for the Yawning Titan autonomous cyber defence environment. We suggest that other, similar or more complex cyber defence simulators could be readily adapted in the same way, due to the fundamental node-based structure of the observation and action spaces. In fact, when designing a simulator, it is likely to be much more straightforward to use an interface such as Entity Gym than devising rules for how a fixed-input agent can interact through the standard Gymnasium interface. One drawback is the lack of out-of-the-box implementations of algorithms designed to work with entity-based environments, and tuned for optimal performance with complex Transformer or GNN policies. Indeed, as can be seen from the training plots, the RogueNet policy displayed occasional instability during training in the environment setting and hyperparameters we used, although this might rectifiable by using a lower learning rate. We believe any shortcomings in out-of-the-box performance of an ‘entity-based’ policy is outweighed by the potential for training and generalisation across variable network topologies, and can be rectified with more development and testing. We suggest that future cyber defence environments be designed using an ‘entity-based’ philosophy, and constructed with support for using an interface akin to Entity Gym.
@@ 8.2 Environment complexity and intra-episode variation
We have demonstrated the ability of the entity-based framework and a Transformer-based policy to train on varying network topologies with constant nodes, as well as to generalise zero-shot to network sizes not seen during training. It is also straightforward to train a policy on varying numbers of nodes during training, where the agent sees a different-sized network with every episode. The agent performs well and also converges to a high reward, although we don’t include this here since the reward function is dependent in-part on the size of the network. In the future, one might expect this kind of variety to be important for training robust agents.
Another important benefit of the Entity Gym, and entity-based or object-oriented RL more generally, is the support for natural variation in the population of entities within an episode, not just between episodes. In a deployment on a real network, the number of nodes observed by a defensive agent may vary over time. This could be due to devices dropping in and out of wireless connection, being manually connected or disconnected by administrators, or communication being blocked by an adversary. In the Yawning Titan environment, it is possible to enable the use of deceptive nodes by the blue agent, which would allow the agent to add new nodes to the environment mid-episode.
@@ 9 Acknowledgments
The authors would like to acknowledge helpful feedback and comments from Liz Bates and Ed Chapman. This work was supported by the Defence and Security Programme at The Alan Turing Institute, funded by the Government Communications Headquarters (GCHQ). Alberto Caron was funded by the Defence Science and Technology Laboratory (DSTL) which is an executive agency of the UK Ministry of Defence providing world class expertise and delivering cutting-edge science and technology for the benefit of the nation and allies.

### v3

@@ 6.2 Generalisation to unseen topologies
The second set of experiments we conducted was aimed at evaluating how well entity-based agents generalise at test-time to network sizes not seen during training. As the baseline MLP policies do not support dynamically sized inputs that are different from the ones used at train-time, they cannot be evaluated on varying network topologies at test-time.
To assess the generalisation ability of entity-based policies across different network sizes, we took three policies trained in the Random training regime, one for each network size of 10, 20 and 40. We then evaluated each of the policies on random networks for each of the network sizes. This means that we evaluated a policy trained on random 10-node networks in the same context of random 10-node networks but also on 20 and 40-node networks (‘out-of-training’ context). All evaluations were carried out over $1000$ episodes and so $1000$ different random networks.
Figure 3 displays the distribution of episodic rewards over 1,000 test episodes in the form of box plots, obtained from the evaluations of the three RogueNet agents trained on different network sizes. Subfigure 3(a) presents evaluations of agents trained on 10, 20 and 40-node networks respectively, and evaluated on random 10-node networks at test-time (eval_rand_k_on_10 with $\mathbf{k\in\{10,20,40\}}$). Subfigures 3(b) and 3(c) present the corresponding evaluations on 20-node networks, 3(b), and 40-node networks, 3(c). Summary statistics of the same experiments are presented in Table 1.
Overall, we can observe a negligible difference in the performance of the three entity-based agents across the different network sizes when evaluated at test-time. For example, the agent trained on 40-node networks performs just as well when evaluated on 10-node networks as the agent that was trained natively on 10-node networks. In fact, it shows slightly lower variance in episodic reward than the agent trained on 10 nodes, with $0.87$ standard deviation against the $1.31$ standard deviation of the native 10-node agent. The only exception is the agent trained on 20-node networks and evaluated on 40-node networks, which shows slightly lower mean reward compared to the other two agents. As can be seen from the summary statistics results in Table 1, this case presents a higher standard deviation in the rewards distribution (0.332) compared to the other two cases (0.187 and 0.161). These results suggest the ability of an entity-based agent to learn a general policy that performs consistently on any network size. The generality of these results would need to be verified in more complex environment settings or simulators, and different varieties of network topologies.
@@ 7 Discussion
In this paper, we have demonstrated the basic functionality and the strong generalisation capabilities of an entity-based RL method through the use of the Entity Gym framework for the Yawning Titan autonomous cyber defence environment. We suggest that other, similar or more complex cyber defence simulators could be readily adapted in the same way, due to the fundamental node-based structure of the observation and action spaces. In fact, when designing a simulator, it is likely to be much more straightforward to use an interface such as Entity Gym than devising rules for how a fixed-input agent can interact through the standard Gymnasium interface. One drawback is the lack of out-of-the-box implementations of algorithms designed to work with entity-based environments, and tuned for optimal performance with complex Transformer or GNN policies. Indeed, as can be seen from the training plots, the RogueNet policy displayed occasional instability during training in the environment setting and hyperparameters we used, although this might be rectifiable by using a lower learning rate. We believe any shortcomings in out-of-the-box performance of an entity-based policy is outweighed by the potential for training and generalisation across variable network topologies, and can be rectified with more development and testing. We suggest that future cyber defence environments be designed using an entity-based philosophy, and constructed with support for using an interface akin to Entity Gym.
@@ 8.2 Environment complexity and intra-episode variation
We have demonstrated the ability of the entity-based framework and a Transformer-based policy to train on varying network topologies with constant nodes, as well as to generalise zero-shot to network sizes not seen during training. It is also straightforward to train a policy on varying numbers of nodes during training, where the agent sees a different-sized network with every episode. The agent performs well and also converges to a high reward, although we do not include this here since the reward function is dependent in-part on the size of the network. In the future, one might expect this kind of variety to be important for training robust agents.
Another important benefit of the Entity Gym, and entity-based or object-oriented RL more generally, is the support for natural variation in the population of entities within an episode, not just between episodes. In a deployment on a real network, the number of nodes observed by a defensive agent may vary over time. This could be due to devices dropping in and out of a wireless connection, being manually connected or disconnected by administrators, or communication being blocked by an adversary. In the Yawning Titan environment, it is possible to enable the use of deceptive nodes by the blue agent, which would allow the agent to add new nodes to the environment mid-episode.
@@ 9 Acknowledgments
The authors would like to acknowledge helpful feedback and comments from Liz Bates and Ed Chapman. This work was supported by the Defence and Security Programme at The Alan Turing Institute, funded by the Government Communications Headquarters (GCHQ). Alberto Caron was funded by the Defence Science and Technology Laboratory (DSTL) which is an executive agency of the UK Ministry of Defence providing world-class expertise and delivering cutting-edge science and technology for the benefit of the nation and allies.